find the coordinates of the foot of the perpendicular from (-2.5) to 3x+y+11=0

Grade:12

2 Answers

Arun
5 years ago
for foot of perpendicular
 
(x + 2)/ 3 =  (y – 5)/ 1 =  – ( – 6 + 5 + 11)/ (9 + 1)
 
now use this and you will find x and y which are the foot of perpendicular
Samyak Jain
5 years ago
Method 1 : Given equation : 3x + y + 11 = 0                …..........(1)
Slope = – 3
Slope of line perpendicular to given line = –1/ –3 = 1/3
\therefore Equation of perpendicular line from (-2.5) to 3x+y+11=0 is
y – 5 = (1/3)(x + 2)   i.e. x – 3y + 17 = 0                      ...............(2)
Point of intersection of lines given by (1) & (2) is (– 5,4)
Method 2 : Parametric form.
Foot of perpendicular from (x, y1) to ax + by + c = 0 can be found by formula 
(x – x1) / a  =  (y – y1) / b  =  (ax1 + by1 + c) / a+ b2
Solving, we get x = – 5 , y = 4
\therefore The required point is (– 5 , 4) .